# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ 4.1 Summary of the Main Results
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+})\right\}.$ (4.2)
That is, the infinitesimal generator ${{\mathbf{G}}}$ of the whole system is the sum of local infinitesimal generators ${{\mathbf{G}}}(x)$. Notably, the local generators commute, i.e., $[{{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]=0$, if $x$ and $x^{\prime}$ are not neighbors. Similarly, the Lagrangian $L(w_{t},w_{t}^{+})$ of the entire system is the sum of the local field Lagrangian $L(w_{t,x},w_{t,x}^{+})$. The transition probability $\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle$ can be formulated in a path integral formalism using the Lagrangian. As we take the limit that the entropy of the system $H_{w}\to 0$, the path that minimizes the integral of the Lagrangian becomes the dominant contribution to the path integral.
@@
${\bar{\gamma}}(x)=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}(t),$ (4.3)
which is the time average of the objective signals it receives over an infinite time horizon. We call an entity to be objective-driven if it evolves to minimize its objective value. To construct such objective-driven dynamics, we may simulate random trials of different local generators and select the most effective one. However, a more intriguing approach would be to enable the system to simulate gradient descent. This leads to the key question: What is the gradient of the local objective value with respect to the local generator? We derive this gradient formula as
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\hat{\bar{\bm{\varphi}}}},\qquad\text{where }\ {{\mathbf{P}}}=1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}.$ (4.4)
@@
${\bm{\Gamma}}_{x}={\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where }\ {\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}.$ (4.5)
But, if the objective propagator is not trivial, the gradient formula would be non-local in general, making it difficult for local entities to estimate its corresponding gradient. We find an interesting class of objective operators
@@ 4.2 Discussion on Different Perspectives
The proposed framework for objective-driven dynamical stochastic fields is derived in a self-contained manner, from three fundamental principles—complete configuration, locality, and purposefulness. These principles are intuitive, echoing concepts found across diverse disciplines. As a result, the framework is potentially interdisciplinary and can be viewed from various perspectives. Given its generality, it connects to a vast body of existing literature. Although a comprehensive overview is beyond our current scope, we acknowledge several major connections. We note, however, that these high-level connections should be viewed as suggestive rather than definitive.
@@
An objective-driven dynamical stochastic field may be conceptualized as a kind of neural network within an environment. From this perspective, the neural network consists of acting entities in the field, while the environment comprises environmental entities. Thus, each acting entity is viewed as a neuron, collectively forming a neural network. These neurons have individual configurations that determine their subsequent actions, including changing its own configuration and interacting with other entities. Similar to how neurons in traditional neural networks transmit signals to connected neurons, entities in this field model exchange signals that influence each other’s behavior. This results in a dynamic and adaptive neural network, where each neuron’s behavior continuously adjusts based on the objective signals it receives from the other entities, including those from the environment. Each neuron also generates and propagates objective signals to other neurons, resembling a feedback propagation mechanism similar to those found in various artificial neural networks.
The prominent model of artificial neural networks (ANNs) are deep neural networks, which are commonly trained via backpropagation (Rumelhart et al., 1986; LeCun et al., 1988; Hecht-Nielsen, 1992). Backpropagation is considered to be non-local in nature, which some argue lacks biological plausibility (Grossberg, 1987; Bengio et al., 2015), and efforts are made to train deep neural networks in more biologically plausible ways (Bengio et al., 2015; Hinton, 2022). Unlike deep neural networks, some ANNs are naturally biologically plausible. Notable examples include models like the Hopfield network (Hopfield, 1982) and Boltzmann machines (Ackley et al., 1985), which are designed for storing information through associative memory. Both of these were influenced by local Hebbian learning theory and are conventionally trained via minimizing certain energy functions. Thus, from the perspective of objective-driven dynamic stochastic fields, this energy function can be viewed as a special case of an objective value, one without dynamics or propagation, that the system aims to minimize. Furthermore, spiking neural networks (Maass, 1997; Gerstner and Kistler, 2002; Yamazaki et al., 2022) offer another layer of biological realism by mimicking the actual electrical activity of neurons, which can be also trained via local Hebbian learning rules.
@@
From a reinforcement learning perspective, each entity in the intelligent field may be viewed as an agent, which makes the setting similar to a partially observable multi-agent RL problem. The objective signals exchanged among entities align with the notion of rewards. However, in intelligent field framework, the notions of “agent” and “environment” are relative and symmetric: from the standpoint of any given agent, the rest of the field (i.e., all other agents) constitutes the environment. Consequently, the system is viewed as a network of interacting agents whose behaviors and objective signals collectively shape one another’s learning processes. Moreover, from this perspective, the reward mechanism is quite complex, as each agent both generates rewards for others and receives rewards from them. There is another important concept in the intelligent field framework that is missing in the RL perspective, i.e., decomposability. In the intelligent field framework, each entity can be composed of smaller and simpler entities, and the complex behavior of a larger entity emerges from the composition of these smaller entities. It thus prompts a unified view of reinforcement learning and neural networks, where an agent is composed of many simpler neuron-agents, each exhibiting basic behaviors. The collective dynamics of these simple neuron-agents give rise to the complex behavior of the larger agent.
@@
Complex systems composed of interacting, dynamic components are ubiquitous in nature. They include examples as diverse as computation models, physical systems, and neural architectures, usually with emergent behaviors that arise from interactions among simpler components (Ottino, 2003; Ladyman et al., 2013). Cellular automata (CA) (Von Neumann et al., 1966), a theoretical model for self-replicating cell grids, are discrete dynamical systems in which each cell updates its state based on local interaction rules. It is shown to be Turing complete (Rendell, 2002; Cook et al., 2004), i.e., given appropriate initial conditions, they can simulate any computable process. The Ising model is another example illustrating how simple local rules (spin alignments) can yield global phase transitions and rich complexity, bridging conceptually to Hopfield networks and Boltzmann machines as discussed preciously. Likewise, random processes on graphs, e.g., epidemic-like spreading, capture how network topology drives complex phenomena in various domains (Easley et al., 2010).
The objective-driven dynamical stochastic field, when the ground space is a discrete graph, can be viewed as another such complex system, in which local node interactions govern overall behavior. For example, from a cellular automata perspective, the intelligent field can be viewed as a generalized framework that extends traditional CA principles by integrating continuous-time stochastic dynamics and self-adaptive learning rules. While it preserves the local-update rule of cellular automata, it departs from static rule sets by introducing stochastic objective signals exchanged among neighboring entities. Each entity then update its internal configuration to minimize the objective values provided by others, guiding the system as a whole toward intended goals.
@@
Our theory begins by introducing a Hilbert space to encode the distributional information, i.e., the “superposition,” of possible system configurations. The dynamics within this Hilbert space are governed by a first-order differential equation, $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}={{\mathbf{G}}}{\hat{\bm{\varphi}}}$, which is a reformulation of the Kolmogorov forward equation. This resembles the Schrödinger’s equation, where the generator of the system dynamics is the Hamiltonian. Moving to a field formulation, we prove that the system’s generator ${{\mathbf{G}}}=\sum_{x}{{\mathbf{G}}}(x)$ is the sum of local generators, where the commutator $[{{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]=0$ for non-neighboring points $x$ and $x^{\prime}$. This mirrors how the total Hamiltonian of a quantum field is obtained as the spatial integral of the Hamiltonian density. Likewise, we can derive the path integral formalism: $\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle=\int{\mathcal{D}}{\omega}\ e^{-S[{\omega}]}$, where $S[{\omega}]=\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})$ is analogous to the action of the path, i.e., the spatial-temporal integral of the Lagrangian density. As we take the limit $H_{w}\to 0$, i.e., letting the entropy of the system goes to zero, the path that minimizes $S[{\omega}]$ dominates the integral, similar to how taking $\hbar\to 0$ in Feynman’s path integral recovers the classical limit in quantum mechanics.
These theoretical tools, inspired by quantum fields, provide a compact model of the dynamical stochastic field, as described in Section 2. From an artificial intelligence perspective, however, it is essential for the model to demonstrate adaptive behavior that is capable of learning to fulfill a specific purpose. Therefore, by further introducing the concept of purposefulness into this system in Section 3, we arrive at the final result of this paper: a theoretical framework for objective-driven dynamical stochastic fields which we refer to as intelligent fields.
We would like to highlight that the connections discussed above serve primarily as illustrative examples to help clarify how intelligent fields relate to various established perspectives. These examples are intended to offer intuitive insights and suggest possible applications. Nevertheless, as mentioned earlier, each of these perspectives differs in certain respects. We recognize that these examples are not comprehensive, and there can exist additional connections that are beyond the current scope. A deeper investigation into these specific links remains an important direction for future work.
@@ 5 Conclusion
In this work, we introduce a theoretical framework for understanding and analyzing objective-driven dynamical stochastic fields which we refer to as intelligent fields. By formalizing the principles of complete configuration, locality, and purposefulness, we establish a mathematical foundation for understanding such intricate systems. In addition, we explore design methodologies from the perspective of artificial intelligence. While the proposed principles and framework provide a foundation for understanding and designing such systems, there remain significant opportunities for further exploration and refinement, ranging from theoretical advancements to practical applications.
@@ Appendix A Proofs in Section 2
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}(t)={{\mathbf{G}}}{\hat{\bm{\varphi}}}(t),$ (A.1)
@@
${\hat{\bm{\varphi}}}(t)=e^{Gt}{\hat{\bm{\varphi}}}(0).$ (A.2)
Note that the exponential of the linear operator is defined as $e^{{{\mathbf{G}}}t}=1+\sum_{n=1}^{\infty}\frac{1}{n!}({{\mathbf{G}}}t)^{n}$. When it is clear in the context, we may simply use ${\hat{\bm{\varphi}}}$ to denote ${\hat{\bm{\varphi}}}(t)$ for notational ease.
@@
Consider a process ${\omega}(t)$ where the probability distribution of its configuration at time $t$ is characterized by ${\hat{\bm{\varphi}}}(t)$. We may omit the time label and simply use ${\hat{\bm{\varphi}}}$ to denote ${\hat{\bm{\varphi}}}(t)$. To verify if any two elements in ${\mathcal{H}}$ are equal, we may verify that they are equal at every basis. Thus, we can prove the proposition by verifying that $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}$ and ${{\mathbf{G}}}{\hat{\bm{\varphi}}}$ are equal at every basis.
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}\rangle$ $=\lim_{{\Delta t}\to 0}\left\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},\frac{{\hat{\bm{\varphi}}}(t+{\Delta t})-{\hat{\bm{\varphi}}}(t)}{{\Delta t}}\right\rangle$ (A.3)
$=\lim_{{\Delta t}\to 0}\frac{1}{{\Delta t}}\left(\left\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{\hat{\bm{\varphi}}}(t+{\Delta t})\right\rangle-\left\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{\hat{\bm{\varphi}}}(t)\right\rangle\right)$ (A.4)
@@
To proceed, we may apply the law of total probability on $p({\omega}(t+{\Delta t})={\omega}^{\prime})$ conditioning on ${\omega}(t)={\omega}$. Moreover, noting that $p({\omega}(t)={\omega})={\varphi}^{{\omega}}$, we have
@@
$=\sum_{\omega}\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle\ {\varphi}^{{\omega}}=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}({\varphi}^{{\omega}}{\hat{\bm{e}}}_{{\omega}})\rangle$ (A.10)
$=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{\varphi}}}\rangle,$ (A.11)
where the last step is by definition (equation 2.4) that ${\hat{\bm{\varphi}}}={\varphi}^{{\omega}}{\hat{\bm{e}}}_{{\omega}}$. Since ${\omega}^{\prime}$ is arbitrary, we must have $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}={{\mathbf{G}}}{\hat{\bm{\varphi}}}$.
This is a standard first order differential equation, where the solution is known to be ${\hat{\bm{\varphi}}}(t)=e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)$. We may verify that
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}(t)=\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)={{\mathbf{G}}}e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)={{\mathbf{G}}}{\hat{\bm{\varphi}}}(t).$ (A.12)
@@
We can construct a linear map ${\mathbf{F}}:{\mathcal{H}}({\Omega}\times{\Omega}^{\prime})\to{\mathcal{H}}({\Omega})\otimes{\mathcal{H}}({\Omega}^{\prime})$ as ${\mathbf{F}}{\hat{\bm{e}}}_{{\omega}{\omega}^{\prime}}={\hat{\bm{e}}}_{{\omega}}\otimes{\hat{\bm{e}}}_{{\omega}^{\prime}}$. By definitions one can verify that it is bijective, and it preserves the inner product. ∎
@@
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\mid\alpha\beta\mu\nu)-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t}),$ (A.15)
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\mu^{\prime}},{{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\mu^{\prime}\mid\alpha\beta\mu)-\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}+o({\Delta t}),$ (A.16)
$\langle{\hat{\bm{e}}}_{\beta^{\prime}\nu^{\prime}},{{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\beta^{\prime}\nu^{\prime}\mid\alpha\beta\nu)-\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}+o({\Delta t}).$ (A.17)

### v2

@@ 4.1 Summary of the Main Results
$\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+})\right\}.$ (4.2)
That is, the infinitesimal generator $\mathbf{G}$}} of the whole system is the sum of local infinitesimal generators ${{\mathbf{G}}}(x)$. Notably, the local generators commute, i.e., ${{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]=0$, if $x$ and $x^{\prime}$ are not neighbors. Similarly, the Lagrangian $L(w_{t},w_{t}^{+})$ of the entire system is the sum of the local field Lagrangian $L(w_{t,x},w_{t,x}^{+})$. The transition probability $\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle$ can be formulated in a path integral formalism using the Lagrangian. As we take the limit that the entropy of the system $H_{w}\to 0$, the path that minimizes the integral of the Lagrangian becomes the dominant contribution to the path integral.
@@
${\bar{\gamma}}(x)=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\varphi\rangle}(t),$ (4.3)
which is the time average of the objective signals it receives over an infinite time horizon. We call an entity objective-driven if it evolves to minimize its objective value. To construct such objective-driven dynamics, we may simulate random trials of different local generators and select the most effective one. However, a more intriguing approach would be to enable the system to simulate gradient descent. This leads to the key question: What is the gradient of the local objective value with respect to the local generator? We derive this gradient formula as
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\lvert\bar{\varphi}\rangle},\qquad\text{where }\ {{\mathbf{P}}}=1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}.$ (4.4)
@@
${\bm{\Gamma}}_{x}=\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where }\ {\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}.$ (4.5)
However, if the objective propagator is not trivial, the gradient formula would be non-local in general, making it difficult for local entities to estimate its corresponding gradient. We find an interesting class of objective operators
@@ 4.2 Discussion on Different Perspectives
The proposed framework for objective-driven dynamical stochastic fields is derived in a self-contained manner, from three fundamental principles: complete configuration, locality, and purposefulness. These principles are intuitive, echoing concepts found across diverse disciplines. As a result, the framework is potentially interdisciplinary and can be viewed from various perspectives. Given its generality, it connects to a vast body of existing literature. Although a comprehensive overview is beyond our current scope, we acknowledge several major connections. We note, however, that these high-level connections should be viewed as suggestive rather than definitive.
@@
An objective-driven dynamical stochastic field may be conceptualized as a kind of neural network within an environment. From this perspective, the neural network consists of acting entities in the field, while the environment comprises environmental entities. Thus, each acting entity is viewed as a neuron, collectively forming a neural network. These neurons have individual configurations that determine their subsequent actions, including changing their own configuration and interacting with other entities. Similar to how neurons in traditional neural networks transmit signals to connected neurons, entities in this field model the exchange of signals that influence each other’s behavior. This results in a dynamic and adaptive neural network, where each neuron’s behavior continuously adjusts based on the objective signals it receives from the other entities, including those from the environment. Each neuron also generates and propagates objective signals to other neurons, resembling a feedback propagation mechanism similar to those found in various artificial neural networks.
The prominent models of artificial neural networks (ANNs) are deep neural networks, which are commonly trained via backpropagation (Rumelhart et al., 1986; LeCun et al., 1988; Hecht-Nielsen, 1992). Backpropagation is considered to be non-local in nature, which some argue lacks biological plausibility (Grossberg, 1987; Bengio et al., 2015), and efforts are made to train deep neural networks in more biologically plausible ways (Bengio et al., 2015; Hinton, 2022). Unlike deep neural networks, some ANNs are naturally biologically plausible. Notable examples include models like the Hopfield network (Hopfield, 1982) and Boltzmann machines (Ackley et al., 1985), which are designed for storing information through associative memory. Both of these were influenced by local Hebbian learning theory and are conventionally trained via minimizing certain energy functions. Thus, from the perspective of objective-driven dynamic stochastic fields, this energy function can be viewed as a special case of an objective value, one without dynamics or propagation, that the system aims to minimize. Furthermore, spiking neural networks (Maass, 1997; Gerstner and Kistler, 2002; Yamazaki et al., 2022) offer another layer of biological realism by mimicking the actual electrical activity of neurons, which can also be trained via local Hebbian learning rules.
@@
From a reinforcement learning perspective, each entity in the intelligent field may be viewed as an agent, which makes the setting similar to a partially observable multi-agent RL problem. The objective signals exchanged among entities align with the notion of rewards. However, in the intelligent field framework, the notions of “agent” and “environment” are relative and symmetric: from the standpoint of any given agent, the rest of the field (i.e., all other agents) constitutes the environment. Consequently, the system is viewed as a network of interacting agents whose behaviors and objective signals collectively shape one another’s learning processes. Moreover, from this perspective, the reward mechanism is quite complex, as each agent both generates rewards for others and receives rewards from them. There is another important concept in the intelligent field framework that is missing from the RL perspective, i.e., decomposability. In the intelligent field framework, each entity can be composed of smaller and simpler entities, and the complex behavior of a larger entity emerges from the composition of these smaller entities. It thus prompts a unified view of reinforcement learning and neural networks, where an agent is composed of many simpler neuron-agents, each exhibiting basic behaviors. The collective dynamics of these simple neuron-agents give rise to the complex behavior of the larger agent.
@@
Complex systems composed of interacting, dynamic components are ubiquitous in nature. They include examples as diverse as computational models, physical systems, and neural architectures, usually with emergent behaviors that arise from interactions among simpler components (Ottino, 2003; Ladyman et al., 2013). Cellular automata (CA) (Von Neumann et al., 1966), a theoretical model for self-replicating cell grids, are discrete dynamical systems in which each cell updates its state based on local interaction rules. It is shown to be Turing complete (Rendell, 2002; Cook et al., 2004), i.e., given appropriate initial conditions, it can simulate any computable process. The Ising model is another example illustrating how simple local rules (spin alignments) can yield global phase transitions and rich complexity, bridging conceptually to Hopfield networks and Boltzmann machines as discussed previously. Likewise, random processes on graphs, e.g., epidemic-like spreading, capture how network topology drives complex phenomena in various domains (Easley et al., 2010).
The objective-driven dynamical stochastic field, when the ground space is a discrete graph, can be viewed as another complex system in which local node interactions govern overall behavior. For example, from a cellular automata perspective, the intelligent field can be viewed as a generalized framework that extends traditional CA principles by integrating continuous-time stochastic dynamics and self-adaptive learning rules. While it preserves the local-update rule of cellular automata, it departs from static rule sets by introducing stochastic objective signals exchanged among neighboring entities. Each entity then updates its internal configuration to minimize the objective values provided by others, guiding the system as a whole toward intended goals.
@@
Our theory begins by introducing a Hilbert space to encode the distributional information, i.e., the “superposition,” of possible system configurations. The dynamics within this Hilbert space are governed by a first-order differential equation, $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\lvert\varphi\rangle}={{\mathbf{G}}}{\lvert\varphi\rangle}$, which is a reformulation of the Kolmogorov forward equation. This resembles the Schrödinger’s equation, where the generator of the system dynamics is the Hamiltonian. Moving to a field formulation, we prove that the system’s generator ${{\mathbf{G}}}=\sum_{x}{{\mathbf{G}}}(x)$ is the sum of local generators, where the commutator ${{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]=0$ for non-neighboring points $x$ and $x^{\prime}$. This mirrors how the total Hamiltonian of a quantum field is obtained as the spatial integral of the Hamiltonian density. Likewise, we can derive the path integral formalism: $\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle=\int{\mathcal{D}}{\omega}\ e^{-S[{\omega}]}$, where $S[{\omega}]=\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})$ is analogous to the action of the path, i.e., the spatial-temporal integral of the Lagrangian density. As we take the limit $H_{w}\to 0$, i.e., letting the entropy of the system go to zero, the path that minimizes $S[{\omega}]$ dominates the integral, similar to how taking $\hbar\to 0$ in Feynman’s path integral recovers the classical limit in quantum mechanics.
These theoretical tools, inspired by quantum fields, provide a compact model of the dynamical stochastic field, as described in Section 2. From an artificial intelligence perspective, however, it is essential for the model to demonstrate adaptive behavior that is capable of learning to fulfill a specific purpose. Therefore, by further introducing the concept of purposefulness into this system in Section 3, we arrive at the final result of this paper: a theoretical framework for objective-driven dynamical stochastic fields, which we refer to as intelligent fields.
We would like to highlight that the connections discussed above serve primarily as illustrative examples to help clarify how intelligent fields relate to various established perspectives. These examples are intended to offer intuitive insights and suggest possible applications. Nevertheless, as mentioned earlier, each of these perspectives differs in certain respects. We recognize that these examples are not comprehensive and that there can be additional connections beyond the current scope. A deeper investigation into these specific links remains an important direction for future work.
@@ 5 Conclusion
In this work, we introduce a theoretical framework for understanding and analyzing objective-driven dynamical stochastic fields, which we refer to as intelligent fields. By formalizing the principles of complete configuration, locality, and purposefulness, we establish a mathematical foundation for understanding such intricate systems. In addition, we explore design methodologies from the perspective of artificial intelligence. While the proposed principles and framework provide a foundation for understanding and designing such systems, there remain significant opportunities for further exploration and refinement, ranging from theoretical advancements to practical applications.
@@ Appendix A Proofs in Section 2
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}\lvert\varphi(t)\rangle={{\mathbf{G}}}\lvert\varphi(t)\rangle,$ (A.1)
@@
$\lvert\varphi(t)\rangle=e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle.$ (A.2)
Note that the exponential of the linear operator is defined as $e^{{{\mathbf{G}}}t}=1+\sum_{n=1}^{\infty}\frac{1}{n!}({{\mathbf{G}}}t)^{n}$. When it is clear in the context, we may simply use $\lvert\varphi\rangle$} to denote $\lvert\varphi(t)\rangle$ for notational ease.
@@
Consider a process ${\omega}(t)$ where the probability distribution of its configuration at time $t$ is characterized by $\lvert\varphi(t)\rangle$. We may omit the time label and simply use $\lvert\varphi\rangle$} to denote $\lvert\varphi(t)\rangle$. To verify if any two elements in $\mathcal{H}$} are equal, we may verify that they are equal in every basis. Thus, we can prove the proposition by verifying that $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\lvert\varphi\rangle}$ and ${{\mathbf{G}}}{\lvert\varphi\rangle}$ are equal in every basis.
$\langle{\omega}^{\prime}\rvert\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\lvert\varphi\rangle}$ $=\lim_{{\Delta t}\to 0}\frac{1}{{\Delta t}}\langle{\omega}^{\prime}\rvert(\lvert\varphi(t+{\Delta t})\rangle-\lvert\varphi(t)\rangle)$ (A.3)
$=\lim_{{\Delta t}\to 0}\frac{1}{{\Delta t}}\left(\langle{\omega}^{\prime}\mid\varphi(t+{\Delta t})\rangle-\langle{\omega}^{\prime}\mid\varphi(t)\rangle\right)$ (A.4)
@@
To proceed, we may apply the law of total probability on $p({\omega}(t+{\Delta t})={\omega}^{\prime})$ conditioning on ${\omega}(t)={\omega}$. Moreover, write $\lvert\varphi\rangle=\sum_{{\omega}}\varphi^{\omega}\lvert{\omega}\rangle$ and we have $p({\omega}(t)={\omega})={\varphi}^{{\omega}}$. Thus,
@@
$=\sum_{\omega}\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle\ {\varphi}^{{\omega}}=\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\left(\sum_{{\omega}}{\varphi}^{{\omega}}\lvert{\omega}\rangle\right)\$(A.10)
$=\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}{\lvert\varphi\rangle}.$ (A.11)
Since $\omega$$^{\prime}$ is arbitrary, we must have $\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\lvert\varphi\rangle}={{\mathbf{G}}}{\lvert\varphi\rangle}$.
This is a standard first order differential equation, where the solution is known to be $\lvert\varphi(t)\rangle=e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle$. We may verify that
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}\lvert\varphi(t)\rangle=\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle={{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle={{\mathbf{G}}}\lvert\varphi(t)\rangle.$ (A.12)
@@
We can construct a linear map ${\mathbf{F}}:{\mathcal{H}}({\Omega}\times{\Omega}^{\prime})\to{\mathcal{H}}({\Omega})\otimes{\mathcal{H}}({\Omega}^{\prime})$ as ${\mathbf{F}}\lvert{\omega}{\omega}^{\prime}\rangle=\lvert{\omega}\rangle\otimes\lvert{\omega}^{\prime}\rangle$. By definitions one can verify that it is bijective, and it preserves the inner product. ∎
@@
$\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert{{\mathbf{G}}}\lvert\alpha\beta\mu\nu\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\mid\alpha\beta\mu\nu)-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t}),$ (A.15)
$\langle\alpha^{\prime}\mu^{\prime}\rvert{{\mathbf{M}}}\lvert\alpha\beta\mu\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\mu^{\prime}\mid\alpha\beta\mu)-\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}+o({\Delta t}),$ (A.16)
$\langle\beta^{\prime}\nu^{\prime}\rvert{{\mathbf{N}}}\lvert\alpha\beta\nu\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\beta^{\prime}\nu^{\prime}\mid\alpha\beta\nu)-\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}+o({\Delta t}).$ (A.17)